​Order: NS 1,500 mL IV at 125​ mL/h for 12 hours.

Since the drop factor is 20​ gtt/mL, the initial flow rate was correctly set at 42​ gtt/min. After six​ hours, 900 mL remain to be infused. The infusion is behind schedule. Compute the new flow rate in​ gtt/min. so that the infusion will finish on time. If flow rates may not be adjusted by more than​ 25% of the original​ rate, to what rate would you reset the drip​ rate, or would you contact the prescriber​ instead?
Answer #1

NS 1500ml IV at 125ml/hr for 12 hours

Drop factor = 20gtt/ml

The flow rate = 1500×20/12×60   =41.6

So the 42 gtt/minute is correct.

After 6 hours, 900 ml fluid remaining.

The new flow rate for the next 6 hours = 900/6 = 150 ml/hour

The new flow rate for the next 6 hours = 900×20/6×60   = 50

                             50 gtt/ minute

The flow rate is not more than 25 % of the initial flow rate. The nurse can set the new rate as 50 gtt/minute.